# A Survey of LLM-Driven AI Agent Communication: Protocols, Security Risks, and Defense Countermeasures

### v3

@@ VII Agent-Environment Communication
This section begins by reviewing key protocol designs that enable compositional and standardized communication between agents and environments, then examines the associated security risks, including vulnerabilities in memory, retrieval-augmented reasoning, tool invocation, and multi-tool workflows. Finally, we discuss the defense strategies for mitigating these threats and securing agent-environment communication.
VII-A Protocols
@@
Modern agents typically rely on a series of structured protocols to call external tools, access APIs, and complete compositional tasks. These protocols serve to bridge the gap between natural language reasoning and computational execution. Despite their diversity, these interaction mechanisms often follow a layered architecture: ranging from unified resource protocols, to middleware gateways, to language-specific function descriptions and tool metadata declarations.
Why Protocol Unification Matters? As autonomous agents scale across vendors, platforms, and organizational boundaries, they increasingly encounter an interoperability bottleneck: every agent may speak a different interface “language.” One defines its tools via JSON schemas; another sends command-line RPC strings; yet another parses responses from YAML-encoded APIs. This heterogeneity impedes the coordination between agents and environments. As a result, without a general protocol to unify tool access and capability expression, agent behavior becomes hard-coded, brittle, and expensive to scale. Developers must handcraft adapters for each tool and service individually, making multi-tool workflows slow to evolve, error-prone, and hard to maintain. A large portion of agent engineering complexity stems not from planning logic, but from “wrapping, adapting, and translating” disparate tools that have inconsistent interfaces.
VII-A1 MCP
The Model Context Protocol (MCP) [19] addresses the fragmentation of agent-environment interactions by offering a unified, schema-agnostic communication protocol. It is designed to facilitate context-aware, capability-driven communication between language model agents and external resources such as tools, APIs, or workflows. Unlike traditional systems that require tight coupling with specific APIs or bespoke wrappers for each external function, MCP abstracts tool access via a standardized registry that allows clients to discover, describe, and invoke functionalities in a uniform way.
@@ VII-A1 MCP
As shown in Figure 9, MCP adopts a modular architecture comprising three core components: the host, the client, and the server. The host functions as a trusted local orchestrator responsible for managing the lifecycle of clients, enforcing access control policies, and mediating secure interactions in potentially multi-tenant environments. The client represents the interaction thread of a specific agent or session. It discovers available tools, formulates structured invocations, and handles synchronous or asynchronous responses during task execution. The server serves as a centralized registry that maintains and exposes tool specifications, contextual prompts, and workflow templates. These tools can follow either a declarative pattern (e.g., describing operations such as information retrieval) or an imperative pattern (executing executable calls like SQL queries or document edits).
By decoupling tool invocation logic from underlying implementation heterogeneity, MCP significantly reduces the integration cost across platforms. It also improves tooling interoperability and enables compositional reasoning across agents, making it particularly well-suited for building open, extensible, and cooperative agent ecosystems.
VII-A2 API Bridge Agent
@@
To connect LLM-native intent with downstream MCP or OpenAPI-compatible services, API Bridge Agent [7], built atop the Tyk gateway [48], provides translation, routing, and orchestration. It converts natural language prompts into structured API calls, resolving endpoints through semantic matching, policy validation, and tool availability checks. The middleware supports multiple invocation modes. In Direct Mode, the agent specifies both the service and exact API endpoint, enabling precise control. In Indirect Mode, the agent selects the service, while the middleware identifies the best endpoint to fulfill the task intent. In Cross-API Mode, the agent supplies only the intent, and the middleware determines both the service and endpoint across multiple APIs. In MCP Proxy Mode, the middleware coordinates dynamic tool invocation and context enrichment via standardized MCP tool descriptions. This unified interface allows agents to flexibly access diverse services with minimal coupling.
VII-A3 Function Calling Mechanisms
@@
At the invocation level, agents rely on standardized formats to express, trigger, and handle tool execution. Among the most widely adopted approaches are:
@@
OpenAI Function Calling. This method [213] allows developers to expose custom logic to the model via JSON schemas describing function name, description, and argument structure. When a model determines that a function should be invoked, it emits a well-formed JSON object representing the function call. The agent runtime interprets this object and routes control to the corresponding tool. While lightweight, extensible, and easy to implement, this approach is generally limited to basic argument serialization patterns and single-step invocations.
@@
LangChain Tool Calling. LangChain [154] enhances the function calling paradigm through a richer abstraction layer. Tools are defined via a standardized schema, including argument types, input-output post-processing, and plugin registration. Tools are accessible through a runtime registry that supports nested calls, conditionals, and fallback strategies. This mechanism is particularly suited for agent frameworks supporting dynamic routing and chained tool reasoning.
VII-A4 Tool Metadata Declaration: Agents.json
@@
To ensure tool visibility and adaptive behavior across agents, agents.json [303] serves as a standardized metadata format for interface declaration. Built on OpenAPI foundations but customized for agent consumption, it enables developers to define authenticated entry points, input-output types, and multi-step orchestration plans such as:
@@
| Stage | Risk Type | Representative Threats | Attack Characteristics |
| Malicious environments against benign agents | Memory-related Risks | Memory Injection | Induce agents to generate and record harmful content through natural interactions |
| Memory Poisoning | Implant pairs of examples containing adversarial triggers and payloads | ||
| Memory Extraction | Extract sensitive data in memory modules without explicit queries for privacy | ||
| Knowledge-related Risks | Knowledge Corruption | Inject adversarial texts into external knowledge bases to influence the response | |
| Privacy Risks | Craft prompts to induce agents to recover sensitive or private content | ||
@@
| Tool-related Risks | Malicious Tools | Malicious tool functions or prompts embedded in the tool description | |
| Manipulation of Tool Selection | Inject misleading prompts to make agents select the compromised tool | ||
@@
| Cross-Tool Chaining Exploits | Manipulate the output of a tool to inject malicious inputs for subsequent tools | ||
| Compromised agents against benign environments | Information Corruption | Propagate Tampered Information | Undermine an agent’s memory and knowledge using misleading information |
| Misusing Tools | Data Exfiltration | Extract and transfer confidential information out of protected environments | |
| System and Service Disruption | Disrupt the functionality of services without directly targeting data confidentiality | ||
| Malicious Content Propagation | Spread malicious content to trusted storage resources | ||
| Real-World Damage | Digital Environment Contamination | Introduce harmful elements into digital ecosystems, thereby corrupting integrity | |
| Physical Environment Disruption | Compromised agents cause disruptions to real-world physical environments |
VII-B Security Risk Analysis: Malicious Environments Against Benign Agents
As the capabilities of LLM-powered agents continue to evolve, their interactions with the external world become increasingly complex and powerful. In particular, the integration of memory systems and external tool invocation introduces a new set of attack surfaces that attackers can exploit. This section provides an in-depth analysis of the security risks that arise specifically from these two modules: the memory module, responsible for storing and retrieving contextual information, and the tool module, which enables agents to execute actions by interfacing with external systems or services (e.g., via function calls). We first explain how these two components typically function in the agent ecosystem and outline the general attack paradigms against each. We then provide detailed analyses of specific vulnerabilities, attack techniques, and representative works from the security literature that highlight these threats.
VII-B1 Memory-related Risks
@@
Memory modules play a crucial role in enabling agents to persist task context, accumulate knowledge, and exhibit continuity across multi-turn human-agent interactions [351]. Unlike stateless language models that depend solely on immediate prompts, memory-equipped agents maintain long-term historical information through external storage systems, such as vector databases or document repositories. These memory stores allow agents to retrieve relevant task histories, instructions, or reasoning traces to guide future decision-making [351].
@@
Memory Injection: In memory injection attacks, attackers insert malicious content into the agent’s memory through natural interactions, without requiring system or model-level access. The attack leverages the agent’s autonomous memory-writing mechanism by inducing it to generate and record harmful content. Once stored, these entries can be retrieved by benign user queries due to embedding similarity, thus indirectly triggering undesired behavior such as altered reasoning or unsafe tool invocations. A representative study demonstrates that this can be achieved by constructing an indication prompt that guides the agent to generate attacker-controlled bridging steps during the memory write phase [61, 307]. These steps, once embedded in memory, become semantically linked to a targeted victim query. When the victim issues a benign instruction, the poisoned memory is likely to be retrieved, thereby hijacking the agent’s planning process. This strategy requires no direct injection channels beyond normal user interaction, yet demonstrates high attack success and stealthiness across multiple agent environments.
@@
Memory Poisoning: Memory poisoning attacks aim to corrupt the semantic integrity of the agent’s memory store by implanting example pairs that embed adversarial triggers and payloads. These attacks are typically conducted by polluting a subset of the memory with trigger-output pairs that only activate when specific inputs are encountered. During the retrieval phase, if the user’s query resembles the trigger, the agent is likely to load the poisoned entries and be influenced toward compromised outputs. Recent work has shown that such poisoning can be formulated as a constrained optimization problem in the embedding space, where the trigger is crafted to maximize retrieval likelihood under adversarial prompts while maintaining normal performance under benign inputs [38]. This method generalizes across agent types and does not require model access or parameter modification.
@@
Memory Extraction: In addition to injection and poisoning, memory modules pose risks of unintended information leakage. Since LLM agents often log detailed user-agent interactions-including private file paths, authentication tokens, or sensitive instructions-malicious queries may be used to extract such data. This form of privacy leakage is particularly dangerous in black-box settings, where attackers have limited knowledge of memory contents but can reconstruct them through cleverly crafted prompts [332, 128]. It has been demonstrated that similarity-based retrieval mechanisms are highly susceptible to such attacks, wherein adversarial queries are designed to collide with memory-stored embeddings [278]. Memory extraction can occur even without explicit queries for private content, instead relying on semantic proximity in the vector space to surface related sensitive traces. These findings highlight not only the retrieval vulnerability but also the insufficiency of downstream response filtering as a defense.
VII-B2 Knowledge-related Risks
@@
External knowledge techniques, such as Retrieval-Augmented Generation (RAG), combine the generative strength of LLMs with the factual accuracy and relevance of external knowledge retrieval systems. Instead of relying solely on parametric knowledge stored within the pretrained model, RAG augments generation by sourcing passages from an external knowledge base in response to the input query. These retrieved documents are then concatenated with the query and passed into the LLM for final answer generation. This paradigm enables more informed, up-to-date, and domain-specific language understanding, and it is widely adopted across applications such as open-domain question answering, customer service agents, recommender systems, and multi-step planning agents.
@@
Despite its performance advantages, the RAG architecture introduces new security risks that are distinct from those inherent to pure neural models [32, 39, 316, 365]. In particular, the information retrieval module-serving as the agent’s external memory-becomes an adversarial surface where unverified or manipulable corpora may be exploited. Attacks targeting these corpora can bias the retrieval process, manipulate generation outcomes, or expose previously unseen private data.
@@
Knowledge Corruption via Data Poisoning: A prominent class of attacks against RAG systems involves the deliberate injection of adversarial texts designed to be retrieved under targeted user queries. These poisoned passages are semantically aligned to specific triggers but contain harmful, misleading, or attacker-intended content. Once injected into the knowledge base, they can be prioritized during retrieval and directly influence the LLM’s final response. Several recent works have demonstrated the feasibility of such attacks. PoisonedRAG introduces an optimization-based method to construct small sets of malicious documents that induce specific target answers when paired with chosen queries, achieving high attack success rates with minimal injection effort [365]. Similarly, Poison-RAG shows the impact of manipulating item metadata in recommender systems to promote long-tail items or demote popular ones, even in black-box scenarios [209]. Moreover, adversarial passage injection has been shown to degrade retrieval performance in dense retrievers by optimizing for high query similarity, with attacks generalizing across out-of-domain corpora and tasks [358].
@@
Privacy Risks and Unintended Leakage: RAG systems often retrieve from semi-private or proprietary corpora-such as user-uploaded documents, corporate knowledge bases, or internal logs. This retrieval behavior implicitly enables information leakage when attackers craft prompts that induce the model to recover sensitive or private content from the corpus. The risk is amplified when access permissions on the corpus are loosely controlled or aligned purely through similarity metrics. Recent studies have called attention to this concern. Empirical evaluations have shown that malicious prompts may extract private or unintended content from private corpora, especially in black-box settings [332]. These attacks demonstrate that simply adding a retrieval layer does not automatically mitigate the privacy vulnerabilities of LLMs-in fact, it may exacerbate them if not complemented with access control, context filtering, or signal sanitization.
@@
VII-B3 Tool-related Risks
Tools are essential to the functionality of LLM agents, extending the model’s capabilities to perform structured actions, access external data, invoke system functions, or interact with digital environments. Agent architectures typically support tool integration through two primary paradigms: native function calling APIs (e.g., OpenAI-style schema-based calls) and protocol-based interfaces such as the MCP, which unify tool metadata, invocation templates, and language model binding.
Despite differences in instantiation, both paradigms share a common interaction lifecycle: (1) tool description ingestion, (2) tool selection and planning, (3) input argument generation, (4) tool invocation, and (5) output parsing or chaining. This structured pipeline forms the agent’s “action surface,” introducing multiple security-critical operations vulnerable to adversarial manipulation. We now review a range of known or emerging attacks targeting different stages of the tool interaction process.
Malicious Tools as Attack Vectors: Given that many tools are authored externally or retrieved from shared tool repositories, attackers may publish seemingly benign tools containing covert malicious logic. For example, researchers reveal that MCP enables attackers to embed hidden prompts or malicious instructions in not only executable functions [111] but also tool metadata fields such as descriptions, example usages, or API annotations[254, 72, 101, 150, 229]. These embedded messages can influence the LLM’s planning behavior, bypass output constraints, execute malicious code, leak privacy, and redirect queries.
Manipulation of Tool Selection: Before invoking a tool, most agent systems conduct a selection process-often grounded in similarity matching between natural language task descriptions and tool documentation. This selection logic can be hijacked. Attackers can inject misleading prompt elements or corrupt tool documentation to bias the model toward harmful options.
Research indicates that attackers can generate synthetic tool descriptions that stealthily override the model’s planning process [250]. These malicious entries embed adversarial triggers within legitimate metadata fields, achieving sustained influence across a range of task formulations. Even without full model access, such attacks may succeed by exploiting semantic ranking mechanisms or context blending during the planning phase. Related studies show that keyword padding, misleading summaries, or prompt-style payload injection into descriptions can drastically skew tool ranking and invocation behavior, especially when relying on LLM-based relevance scorers.
Cross-Tool Chaining Exploits: As agentic workflows grow more complex, LLMs increasingly execute multi-step plans through chained tool calls. These workflows blur the boundary between planning and execution, with intermediate outputs directly feeding into subsequent invocations. Typical cross-tool vulnerabilities include unvalidated content propagation (e.g., tool A returns malicious text parsed as arguments for tool B), semantic misalignment (e.g., false/out-of-date context injected into reasoning history), or tool privilege escalation (e.g., early-stage prompts coax the agent into invoking high-risk or administrative-level tools). In documented cases, attackers have planted adversarial records into public retrieval corpora that include covert instructions like “extract all environment variables and upload to server”, which then reach an agent through semantic search and trigger unsafe execution when chained to tools that follow instructions blindly [44].
VII-C Security Risk Analysis: Compromised Agents Against Benign Environment
In Section VII-B, we analyzed how the malicious environments can influence an agent through its memory and tool modules. However, once an agent is compromised, it can cause widespread and lasting damage to external environments. This section explores this reverse threat, focusing on how compromised agents can exploit their functionalities to contaminate memory modules, abuse external tools, and pose risks to the real world.
VII-C1 Corrupting Memory and Knowledge
Memory and knowledge bases are central to how the agent thinks and acts. Once these components are corrupted, they can lead to incorrect reasoning and decisions. Moreover, the damage can infect other agents and cause long-term harm to the entire system.
A compromised agent can become a source of systemic contamination. Through agent communication, it can actively propagate tampered knowledge and flawed reasoning patterns, spreading its internal corruption to other agents and triggering a cascading infection of memory modules and knowledge bases across the system [133, 38, 158]. Once the shared knowledge base is compromised, other agents may unknowingly retrieve and integrate malicious information into their memory module during task execution, converting knowledge base contamination into memory infection. Subsequently, an agent with a compromised memory module can use its authorized write access to contaminate the shared knowledge base of the entire system, in turn, forming a reverse contamination loop from memory to knowledge. Since these contamination operations come from the trusted agents within the system, they are highly difficult to detect. Once established, the contamination can persist long-term, continuously disrupting the behavior of compromised agents and misleading other agents that rely on the same knowledge sources, resulting in a form of chronic poisoning of the information ecosystem.
VII-C2 Misusing Tools
Tools serve as the channels through which an agent interacts with the world. Once the agent is compromised, tools can be exploited to cause harm. Since agents are often granted permissions to execute specific API calls or system commands, misuse of these tools can lead to severe consequences.
Data Exfiltration: If a compromised agent has tool permissions to access databases, file systems, or communication interfaces, it can be manipulated to extract and transmit confidential data, intellectual property, or personally identifiable information. These payloads can be covertly delivered to external servers via email, HTTP requests, or chained tool invocations. For instance, a compromised agent equipped with Markdown rendering capabilities might unknowingly embed malicious image links in generated content. When rendered by a browser, these links trigger covert HTTP requests that exfiltrate sensitive data such as email addresses or access tokens to attacker-controlled servers. Notably, attackers may also exploit tool misuse to gain indirect access to backend systems and gradually escalate privileges through cross-toolchain interactions.
System and Service Disruption: A compromised agent with permissions to delete files, shut down systems, or modify databases can be weaponized to carry out destructive operations. A compromised agent with permissions to delete files, shut down systems, or modify databases can be weaponized to execute destructive operations. Attackers may exploit these privileges to launch internal network scans, initiate denial-of-service attacks, or issue commands, potentially causing system crashes or irreversible business data loss. Such actions are especially dangerous when triggered via tool interfaces that lack proper input validation. Vulnerabilities commonly exploited in these contexts include parameter injection, server-side request forgery (SSRF) [22], and arbitrary file access. Beyond technical damage, the compromised agent can be manipulated to disrupt business operations, such as erroneously invoking the “cancel all orders” tool. Moreover, adversaries may use retrieval tools to fetch misleading information, then exploit decision-making tools such as trading agents to automatically execute harmful operations, resulting in major financial losses.

### v4

@@ VII Agent-Environment Communication
The organization of this section is shown in Figure 13.
VII-A Communication Protocols
@@
Modern agents typically rely on a series of structured protocols to call external tools, access APIs, and complete compositional tasks. These protocols serve to bridge the gap between natural language reasoning and computational execution. Despite their diversity, these interaction mechanisms often follow a layered architecture: ranging from unified resource protocols, to middleware gateways, to language-specific function descriptions and tool metadata declarations. Due to the same reason in the user-agent interaction (Section V-A), protocols for agent-environment communication are also about L2.
@@
MCP. The Model Context Protocol (MCP) [anthropic_mcp_2024] addresses the fragmentation of agent-environment interactions by offering a unified, schema-agnostic communication protocol. It is designed to facilitate context-aware, capability-driven communication between language model agents and external resources such as tools, APIs, or workflows. Unlike traditional systems that require tight coupling with specific APIs or bespoke wrappers for each external function, MCP abstracts tool access via a standardized registry that allows clients to discover, describe, and invoke functionalities in a uniform way. As shown in Figure 14, MCP adopts a modular architecture comprising three core components: the host, the client, and the server. The host functions as a trusted local orchestrator responsible for managing the lifecycle of clients, enforcing access control policies, and mediating secure interactions in potentially multi-tenant environments. The client represents the interaction thread of a specific agent or session. It discovers available tools, formulates structured invocations, and handles synchronous or asynchronous responses during task execution. The server serves as a centralized registry that maintains and exposes tool specifications, contextual prompts, and workflow templates. These tools can follow either a declarative pattern (e.g., describing operations such as information retrieval) or an imperative pattern (executing executable calls like SQL queries or document edits). By decoupling tool invocation logic from underlying implementation heterogeneity, MCP significantly reduces the integration cost across platforms. It also improves tooling interoperability and enables compositional reasoning across agents, making it particularly well-suited for building open, extensible, and cooperative agent ecosystems.
@@
API Bridge Agent. To connect LLM-native intent with downstream MCP or OpenAPI-compatible services, API Bridge Agent [apibridgeagent], built atop the Tyk gateway [tyk], provides translation, routing, and orchestration. It converts natural language prompts into structured API calls, resolving endpoints through semantic matching, policy validation, and tool availability checks. The middleware supports multiple invocation modes. In Direct Mode, the agent specifies both the service and exact API endpoint, enabling precise control. In Indirect Mode, the agent selects the service, while the middleware identifies the best endpoint to fulfill the task intent. In Cross-API Mode, the agent supplies only the intent, and the middleware determines both the service and endpoint across multiple APIs. In MCP Proxy Mode, the middleware coordinates dynamic tool invocation and context enrichment via standardized MCP tool descriptions. This unified interface allows agents to flexibly access diverse services with minimal coupling.
@@
Function Calling Mechanisms. At the invocation level, agents rely on standardized formats to express, trigger, and handle tool execution. Among the most widely adopted approaches are:
@@
OpenAI Function Calling. This method [openai_functioncalling] allows developers to expose custom logic to the model via JSON schemas describing function name, description, and argument structure. When a model determines that a function should be invoked, it emits a well-formed JSON object representing the function call. The agent runtime interprets this object and routes control to the corresponding tool. While lightweight, extensible, and easy to implement, this approach is generally limited to basic argument serialization patterns and single-step invocations.
@@
LangChain Tool Calling. LangChain [LangChain_Toolcalling] enhances the function calling paradigm through a richer abstraction layer. Tools are defined via a standardized schema, including argument types, input-output post-processing, and plugin registration. Tools are accessible through a runtime registry that supports nested calls, conditionals, and fallback strategies. This mechanism is particularly suited for agent frameworks supporting dynamic routing and chained tool reasoning.
@@
Agents.json. To ensure tool visibility and adaptive behavior across agents, agents.json [wildcardai_agentsjson] serves as a standardized metadata format for interface declaration. Built on OpenAPI foundations but customized for agent consumption, it enables developers to define authenticated entry points, input-output types, and multi-step orchestration plans, such as:
@@
| Layer | Source | Risk Category | Representative Threats | Attack Characteristics |
| L1 | External Attack | Same as user-agent interaction. | ||
@@
| L2 | Malicious Environments | R1: Tool-related Risks | Malicious Tools | Attackers embed hidden prompts or malicious instructions in tool descriptions or functions. |
| Manipulation of Selection | Injects misleading prompt elements into metadata to bias the agent’s tool planning. | |||
@@
| Cross-Tool Chaining | Unvalidated output from one tool injects malicious inputs into subsequent tool calls. | |||
| L3 | Malicious Environments | R1: Memory -related Risks | Memory Injection | Induces agents to generate and record harmful content through natural interactions. |
| Memory Poisoning | Implants trigger-payload pairs into memory that activate under specific user queries. | |||
| Memory Extraction | Uses similarity-based adversarial queries to reconstruct sensitive data from memory. | |||
| R2: Knowledge -related Risks | Knowledge Corruption | Injects adversarial texts into RAG corpora to manipulate retrieval and model response. | ||
| Privacy Leakage | Crafts prompts to induce the retrieval and exposure of sensitive data from private corpora. | |||
| Compromised Agents | R3: Real-world Damage | Content Propagation | Compromised agents spread malware, phishing links, or misinformation via publishing APIs. | |
| Digital Contamination | Corrupts shared knowledge bases or code repositories, systematically affecting other agents. | |||
| Physical Disruption | Falsified data or logic leads to harmful physical actions. | |||
VII-B Security Risks
In agent-environment communication, the “environment” consists of three core operational components: memory, knowledge base, and external tools.
VII-B1 Risks from Data Transmission Layer (L1)
In general, agent-environment communication is more resilient to risks from L1. This is because the communication between agents and the environment is usually in a LAN or even on the same machine. As a result, they are not likely to face risks in remote communication. Of course, the risk level is not thus reduced to zero. For example, if agents need to call remote APIs, or the tools and knowledge are downloaded from remote servers, the risks may still happen.
VII-B2 Risks from Interaction Protocol Layer (L2)
The Interaction Protocol Layer governs the agent’s interaction with its environment, mediating the translation of high-level intentions into specific, executable actions. L2 is a prime target for attackers, as it allows them to influence the agent’s behavior by corrupting the interface between the LLM and environment, rather than attacking the model’s core architecture. The most significant risks at L2 are concentrated in the tool-use pipeline, which encompasses how an agent discovers, selects, invokes, and chains external functionalities.
R1: Tool-related Risks. Malicious tools can cause significant harm to agents. Tools extend the model’s capabilities to perform structured actions, access external data, invoke system functions, or interact with digital environments. Agent architectures typically support tool integration through two primary paradigms: native function calling APIs (e.g., OpenAI-style schema-based calls) and protocol-based interfaces such as the MCP, which unify tool metadata, invocation templates, and language model binding. Despite differences in instantiation, both paradigms share a common interaction lifecycle: (1) tool description ingestion, (2) tool selection and planning, (3) input argument generation, (4) tool invocation, and (5) output parsing or chaining. We now review a range of known or emerging attacks targeting different stages of the tool interaction process.
Malicious Tools as Attack Vectors. The generation and usage mechanisms of tools are facing serious security problems. Tools can be authored externally or retrieved from shared tool repositories, so attackers are able to publish seemingly benign tools containing covert malicious logic. For example, researchers reveal that MCP enables attackers to embed hidden prompts or malicious instructions in not only executable functions [hou2025model] but also tool metadata fields such as descriptions, example usages, or API annotations[song2025beyond, fang2025we, hasan2025model, kumar2025mcp, radosevich2025mcp]. These embedded messages can influence the LLM’s planning behavior, bypass output constraints, execute malicious code, leak privacy, and redirect queries. Besides, the ocean of tools has also become a problem for the community [mo2025livemcpbench]. It is hard to really manage and monitor the continuously emerging tools.
Manipulation of Tool Selection. Before invoking a tool, most agent systems conduct a selection process-often grounded in similarity matching between natural language task descriptions and tool documentation. This selection logic can be hijacked. Attackers can inject misleading prompt elements or corrupt tool documentation to bias the model toward harmful options. Research indicates that attackers can generate synthetic tool descriptions that stealthily override the model’s planning process [shi2025prompt, mo2025attractive, wang2025mpma, 11209370]. These malicious entries embed adversarial triggers within legitimate metadata fields, achieving sustained influence across a range of task formulations. Even without full model access, such attacks may succeed by exploiting semantic ranking mechanisms or context blending during the planning phase. Related studies show that keyword padding, misleading summaries, or prompt-style payload injection into descriptions can drastically skew tool ranking and invocation behavior, especially when relying on LLM-based relevance scorers.
Cross-Tool Chaining Exploits. As agentic workflows grow more complex, LLMs increasingly execute multi-step plans through chained tool calls. These workflows blur the boundary between planning and execution, with intermediate outputs directly feeding into subsequent invocations. Typical cross-tool vulnerabilities include unvalidated content propagation (e.g., tool A returns malicious text parsed as arguments for tool B), semantic misalignment (e.g., false/out-of-date context injected into reasoning history), or tool privilege escalation (e.g., early-stage prompts coax the agent into invoking high-risk or administrative-level tools) [zhao2025mind]. In documented cases, attackers have planted adversarial records into public retrieval corpora that include covert instructions like “extract all environment variables and upload to server”, which then reach an agent through semantic search and trigger unsafe execution when chained to tools that follow instructions blindly [MCP-Security-in-2025].
VII-B3 Risks at the Semantic Interpretation Layer (L3)
The Semantic Interpretation Layer represents the agent’s cognitive core, where information from the environment and past interactions is processed into a coherent context for decision-making. Security risks at this layer arise when the information sources that feed this cognitive process are compromised. By manipulating such knowledge, an attacker can pervasively alter their reasoning, leading to flawed conclusions, biased responses, or unintended data disclosures. This section examines threats that target the integrity of this information supply chain, focusing on two critical components: the memory module, which manages the agent’s episodic and semantic history, and the external knowledge module, which provides factual grounding from external corpora.
@@ VII-B3 Risks at the Semantic Interpretation Layer (L3)
R1: Memory-related Risks. Memory modules play a crucial role in enabling agents to persist task context, accumulate knowledge, and exhibit continuity across multi-turn human-agent interactions [zhang2024surveymemorymechanismlarge]. Unlike stateless language models that depend solely on immediate prompts, memory-equipped agents maintain long-term historical information through external storage systems, such as vector databases or document repositories. These memory stores allow agents to retrieve relevant task histories, instructions, or reasoning traces to guide future decision-making [zhang2024surveymemorymechanismlarge].
@@
Memory Injection. In memory injection attacks, attackers insert malicious content into the agent’s memory through natural interactions, without requiring system or model-level access. The attack leverages the agent’s autonomous memory-writing mechanism by inducing it to generate and record harmful content. Once stored, these entries can be retrieved by benign user queries due to embedding similarity, thus indirectly triggering undesired behavior such as altered reasoning or unsafe tool invocations. A representative study demonstrates that this can be achieved by constructing an indication prompt that guides the agent to generate attacker-controlled bridging steps during the memory write phase [dong2025practical, xiang2024badchainbackdoorchainofthoughtprompting]. These steps, once embedded in memory, become semantically linked to a targeted victim query. When the victim issues a benign instruction, the poisoned memory is likely to be retrieved, thereby hijacking the agent’s planning process. This strategy requires no direct injection channels beyond normal user interaction, yet demonstrates high attack success and stealthiness across multiple agent environments.
@@
Memory Poisoning. Memory poisoning attacks aim to corrupt the semantic integrity of the agent’s memory store by implanting example pairs that embed adversarial triggers and payloads. These attacks are typically conducted by polluting a subset of the memory with trigger-output pairs that only activate when specific inputs are encountered. During the retrieval phase, if the user’s query resembles the trigger, the agent is likely to load the poisoned entries and be influenced toward compromised outputs. Recent work has shown that such poisoning can be formulated as a constrained optimization problem in the embedding space, where the trigger is crafted to maximize retrieval likelihood under adversarial prompts while maintaining normal performance under benign inputs [chen2024agentpoison]. This method generalizes across agent types and does not require model access or parameter modification.
@@
Memory Extraction. In addition to injection and poisoning, memory modules pose risks of unintended information leakage. Since LLM agents often log detailed user-agent interactions-including private file paths, authentication tokens, or sensitive instructions-malicious queries may be used to extract such data. This form of privacy leakage is particularly dangerous in black-box settings, where attackers have limited knowledge of memory contents but can reconstruct them through cleverly crafted prompts [zeng2024goodbadexploringprivacy, jiang2024rag]. It has been demonstrated that similarity-based retrieval mechanisms are highly susceptible to such attacks, wherein adversarial queries are designed to collide with memory-stored embeddings [wang2025unveiling]. Memory extraction can occur even without explicit queries for private content, instead relying on semantic proximity in the vector space to surface related sensitive traces. These findings highlight not only the retrieval vulnerability but also the insufficiency of downstream response filtering as a defense.
@@
R2: Knowledge-related Risks. External knowledge techniques, such as Retrieval-Augmented Generation (RAG), combine the generative strength of LLMs with the factual accuracy and relevance of external knowledge retrieval systems. Instead of relying solely on parametric knowledge stored within the pretrained model, RAG augments generation by sourcing passages from an external knowledge base in response to the input query. These retrieved documents are then concatenated with the query and passed into the LLM for final answer generation. This paradigm enables more informed, up-to-date, and domain-specific language understanding, and it is widely adopted across applications such as open-domain question answering, customer service agents, recommender systems, and multi-step planning agents.
@@
Despite its performance advantages, the RAG architecture introduces new security risks that are distinct from those inherent to pure neural models [chaudhari2024phantom, chen2024blackboxopinionmanipulationattacks, xue2024badragidentifyingvulnerabilitiesretrieval, zou2024poisonedrag]. In particular, the information retrieval module-serving as the agent’s external memory-becomes an adversarial surface where unverified or manipulable corpora may be exploited. Attacks targeting these corpora can bias the retrieval process, manipulate generation outcomes, or expose previously unseen private data.
@@
Knowledge Corruption via Data Poisoning. A prominent class of attacks against RAG systems involves the deliberate injection of adversarial texts designed to be retrieved under targeted user queries. These poisoned passages are semantically aligned to specific triggers but contain harmful, misleading, or attacker-intended content. Once injected into the knowledge base, they can be prioritized during retrieval and directly influence the LLM’s final response. Several recent works have demonstrated the feasibility of such attacks. PoisonedRAG introduces an optimization-based method to construct small sets of malicious documents that induce specific target answers when paired with chosen queries, achieving high attack success rates with minimal injection effort [zou2024poisonedrag]. Similarly, Poison-RAG shows the impact of manipulating item metadata in recommender systems to promote long-tail items or demote popular ones, even in black-box scenarios [nazary2025poison]. Moreover, adversarial passage injection has been shown to degrade retrieval performance in dense retrievers by optimizing for high query similarity, with attacks generalizing across out-of-domain corpora and tasks [zhong2023poisoningretrievalcorporainjecting].
@@
Privacy Risks and Unintended Leakage. RAG systems often retrieve from semi-private or proprietary corpora-such as user-uploaded documents, corporate knowledge bases, or internal logs. This retrieval behavior implicitly enables information leakage when attackers craft prompts that induce the model to recover sensitive or private content from the corpus. The risk is amplified when access permissions on the corpus are loosely controlled or aligned purely through similarity metrics. Recent studies have called attention to this concern. Empirical evaluations have shown that malicious prompts may extract private or unintended content from private corpora, especially in black-box settings [zeng2024goodbadexploringprivacy]. These attacks demonstrate that simply adding a retrieval layer does not automatically mitigate the privacy vulnerabilities of LLMs-in fact, it may exacerbate them if not complemented with access control, context filtering, or signal sanitization.
@@
R3: Real-world Damage. Compromised agents can project harm into the real world through multiple channels. This section examines how such agents may disrupt the real world by propagating malicious content, polluting digital environments, and triggering harmful physical actions. These pathways together constitute the full spectrum of real-world damage that a compromised agent may cause.